OpenStudy (anonymous):

If, a=log(base24)12, b=log(base36)24, c=log(base48)36, then 1+abc=? a) 2ab c)2bc b) 2ac d) 0

OpenStudy (anonymous):

@nitz

OpenStudy (anonymous):

\[a=\log_{24}12\]\[b=\log_{36}24\]\[c=\log_{48}36\]

OpenStudy (anonymous):

this is similar to previous problem

OpenStudy (anonymous):

still, help.

OpenStudy (anonymous):

abc=log(base 48) 12

OpenStudy (anonymous):

i am getting this

OpenStudy (anonymous):

me too got the same.

OpenStudy (shubhamsrg):

maybe change every base to 12 i.e. a=1/log24 ,,b = log24/log36 ,c= log36/log48 =>abc = 1/log48 1+ abc = 1 + 1/log48 = 1/log12 + 1/log48 = log(48*12)/log48 = log(12^2) + log(4) / log48 = 2 + 2log2 /log48 = 2(1 + log 2)/log48 = 2( log 24)/log48 = 2bc phew!! :P

OpenStudy (anonymous):

cant say but still i am trying

OpenStudy (anonymous):

did you get shubhamsrg method???????

OpenStudy (anonymous):

wait.

OpenStudy (anonymous):

@shubhamsrg "1 + 1/log48 = 1/log12 + 1/log48 " . How?

OpenStudy (shubhamsrg):

as log12 = 1 recall that base is 12 everywhere..

OpenStudy (anonymous):

ok thanks. I got it.